kaylene is 30 yrs. old and jessica is 22 yrs. old. how many years ago was kaylene twice as old as jessica?
Answer by rajagopalan(174)

Let the number of years =x
then x years ago....
kaylene was 30 - x
jessica was 22-x
But according to condition, 2*(22-x)=30-x
44-2x=30-x
44-30=-x+2x
14=x
So the answer is 14 years
Verify:
14 years ago.... kaylene was 30 - x =30-14=16
14 years ago.... jessica was 22-x =22-14=8
then 16/8 = 2 ok